The equilibrium constant, Kp, is 4.51 x 10-35 at 450c for the reaction expressed below. N2 (g) + 3H2 (g) equilibrium arrows 2 NH3 (g)

a. Write the equilibrium expression for, Kp, for the reaction.

b. Supposed you start out with only reactants in a rigid container. The initial partial pressure of N2 (g) is 1 atm and that of H2 is 1.4 atm. What ate the partial pressure of each species when the system reaches equilibrium?

Can you explain step by step, please?
Answer #1

A) Kp =(P​​​​​​NH3)2/(P​​​​​​H2​)3(P​N2)

b)

N2 3H2-----> 2NH3
I 1 1.4 0
C -X -3X +2X
E 1-X 1.4-3X 2X

Kp =4.51×10-35 =

(2X)2/(1.4-3X)3(1-X)

As Kp is very very small, so we can neglect X in compersion to 1.4 & 1

So,

4X2/1.4 =4.51×10-35

X = 3.97×10-18 atm

Partial pressure of N2 = 1.0 atm

Partial pressure of H2=1.4atm

Partial pressure of NH3 =  

= 7.94×10-18 atm